OpenStudy (anonymous):

Suppose that you are the assistant to an ecologist who is interested in the formic acid content of ants. Your job is to determine the formic acid content of the ants. You have ground up an ant with a mass of 1.3248 grams. Water was then added and the solution was titrated with a 29.84 ml of 0.2566 M NaOH. Formic acid and sodium hydroxide react with a one to one mole ratio. The molar mass of formic acid is 46.03 g/mol. What is the mass percent of the formic acid in the ant? (mass of acid/mass of ant)

OpenStudy (anonymous):

First find moles of NaOH \[.2566*.02984=.007657 mol NOH\] One to one ratio so \[.007657 mol NaOH= .007657 mol Formic acid\]

OpenStudy (anonymous):

Multiply by the molar mass \[.007657*46.03g/mol = .3525g\]

OpenStudy (anonymous):

Then the mass percent can be found \[.3525/1.3248= .2661*100= 26.61%\]

OpenStudy (anonymous):

26.61%

OpenStudy (anonymous):

The moles were found through the molarity equation \[M=mol/L \] so \[mol=M*L\]

OpenStudy (anonymous):

THANK YOU SO MUCH!! I really appreciate it